# A glance to Luttinger liquid and its platforms

### v1

@@ II.1 Hamiltonian and observables of the Tomonaga-Luttinger liquid
with $[\phi(x),\theta(x^{\prime})]=i\pi\Theta_{H}(x^{\prime}-x)$, $\Theta_{H}$ being the Heaviside step function, while $u$ is the velocity of low-energy excitations and $K$ is the dimensionless Tomonaga-Luttinger liquid parameter. The fermion density is expressed in terms of the operator $\phi(x)$ as
@@
for the boson annihilation operator, while (2) gives the boson density. More generally, bosons with repulsive interactions are in a Tomonaga-Luttinger liquid phase[66]. Independently of particle statistics, the charge stiffness[101] is $\mathcal{D}=uK$ while the compressibility is $\chi=K/(\pi\rho_{0}^{2}u)$, and at low temperature, the specific heat is $C_{v}=\frac{\pi k_{B}^{2}T}{3u}$. These relations allow a non-perturbative definition of $u$ and $K$. Moreover, since hard core bosons can be mapped to spin-1/2 operators[114], with $S_{j}^{+}\sim(-)^{j}\psi_{B}^{\dagger}(ja)$ and $S_{j}^{z}=\rho(x)-1/(2a)$, XXZ spin-1/2 chains also host a Tomonaga-Luttinger liquid phase.[109, 65] A fully polarized spin state maps to empty or full Jordan-Wigner fermion band, and a state with partial magnetixation corresponds to the partial filling of the fermion band and is a TLL. The role of the compressibility is played by the magnetic susceptibility, $\chi=(g\mu_{B})^{2}K/(\pi u)$ where $\mu_{B}$ is the Bohr magneton and $g$ the $g$-factor, while charge stiffness is replaced by spin stiffness[142]. The magnetic heat capacity $C_{m}=\pi k_{B}^{2}T/(3u)$ has a linear temperature dependence, and the Wilson ratio, $R_{W}=\frac{4}{3}\left(\frac{\pi k_{B}}{g\mu_{B}}\right)^{2}\frac{\chi T}{C_{m}}=4K$ measures the Tomonaga-Luttinger exponent[117]. In the case of spin-1/2 fermions, the two-component Tomonaga-Luttinger liquid Hamiltonian reads
@@
with $\psi_{\sigma}(x)=\sum_{r}e^{ir\pi\rho_{0}x/2}\psi_{r,\sigma}(x)$. At half filling, $\rho_{0}=1/a$ replacing the operator $\phi_{c}$ with 00 in the spin density (8 gives the bosonized expression of the spin operators in the spin-1/2 Heisenberg chain. In the helical liquid[169, 172], spin and Fermi momentum are locked together, making the Tomonaga-Luttinger liquid single component [Eq. (1)] as in the case of spinless fermions. The fermion annihilation operators reduce to
@@ II.2 Ground state correlation functions
where $\alpha$ is a short distance cutoff. The power-law singularities in Eq. (14) are determined by $n,m$ and the Tomonaga-Luttinger liquid exponent $K$. From the correlation functions above, one finds the generalized susceptibilities $\chi_{n,m}(q,\omega)$ at pulsation $\omega$ and wavevector $q$ using Fourier transformation. They present power law singularities[56]
$\chi_{n,m}(q,\omega)\sim\left[\frac{(\omega-uq)\alpha}{u}\right]^{\frac{1}{4}\left(\frac{n}{\sqrt{K}}+m\sqrt{K}\right)^{2}-1}\left[\frac{(\omega+uq)\alpha}{u}\right]^{\frac{1}{4}\left(\frac{n}{\sqrt{K}}-m\sqrt{K}\right)^{2}-1},$ (15)
@@ III Itinerant condensed matter TLLs
Itinerant condensed matter systems, such as Tomonaga-Luttinger liquids (TLLs), represent a fascinating class of one-dimensional quantum systems characterized by their distinct properties from traditional Fermi liquids. Unlike conventional three-dimensional metals, where electron interactions can often be treated as perturbative, TLLs exhibit strong correlations that fundamentally alter their behavior. The theoretical framework of TLL provides a robust description of these phenomena. It applies not only to quantum wires[150] and carbon nanotubes[16, 180] but also to edge states[167] in the fractional quantum Hall effect and certain quasi-one-dimensional[164, 83] conductors, highlighting the universal nature of these one-dimensional systems in condensed matter physics. One specific example of itinerant condensed matter TLL is the 1D array of Josephson junctions.
@@ III.1 Josephson junction chains
In the absence of quantum fluctuations, the system would be analogous to the classical pinning of an elastic chain.[53] Such a chain is characterized by a localization length known as the Larkin length, $N_{\mathrm{L}}$. Classically, one would need to apply a force, here an external voltage, to initiate a flow of current. For chains larger than $N_{\mathrm{L}}$, the pinning force is the critical voltage divided by the number of junctions in the chain, and has dependence, $eV_{\mathrm{c}}/N\propto N_{\mathrm{L}}^{-2}\propto E_{\mathrm{0}}^{-1/3}W^{4/3}$, where $E_{\mathrm{0}}$ is the elastic energy, and W is he width of the disorder, equal to the Bloch bandwidth of the junction energy band for maximal disorder. We note $W\sim\hbar\omega_{p}$, a characteristic energy of the junctions. In order to experimentally compare chain families of widely varying $\hbar\omega_{p}$, and chain length $N$, we express the critical voltage and Bloch bandwidth as dimensionless variables, $v\equiv eV_{\mathrm{c}}/N\hbar\omega_{p}$, and $w\equiv W/\hbar\omega_{p}$. In the classical limit, $v=aw^{4/3}$, with prefactor $a=b\left(K/\Lambda\right)^{1/3}$, $b$ a constant $\mathcal{O}\left(1\right)$, and $\Lambda$ the electrostatic screening length.
@@ V Edge states and hinge states of topological matter
A natural realization of 1D electronic structure arises within the boundary modes of 2D topological insulators (TIs) (Fig. 3). Different from conventional electrical insulators (or semiconductor) in which conduction and valence bands are strictly separated by the electronic band gap, topological insulators possess a band-gap only in the bulk of the crystal while metallic boundary states persist on their surfaces and edges. Regardless of dimensionality, what makes topological boundary modes fundamentally different from their trivial counterparts (such as e.g. Shockley [143] or Tamm[148] states) is that (i) their existence is ensured by topological invariants of the bulk band structure – a principle called bulk-boundary correspondence [70] – and that (ii) these systems often carry (at least approximately) linearly dispersing 1D Dirac fermions at their boundaries that are either chiral or helical, spin-degenerate or spinless, depending on which class of topological insulator is being considered (Fig. 3d).
@@
The first-known 2D topological phase of matter is the quantum Hall effect (QHE) [99], later put into the context of band topology by Thouless [155]. In the QHE, a strong magnetic field $H$ externally applied perpendicular to the plane of a 2D electron gas forces electrons into closed trajectories leading to the opening of a gap in the 2D bulk density of states (Landau quantization). Gapless chiral edge states[166] persist at the boundary, spin-degenerate at low magnetic field (Fig. 3b).
@@ V.2 Experimental demonstrations of chiral and helical TLLs
Luttinger liquids have so far been only been reported for the chiral boundaries of fractional quantum Hall states [32, 31, 72, 126, 163, 35] and within the helical edges of QSH insulators [107, 146, 84, 145] insulators. No Luttinger liquid claims have been made, to our knowledge, for the chiral edges of QAH (or Chern) insulators [175, 33] so far.
@@
Experimental probes of both chiral and helical Luttinger liquids in 2D TIs proceed identically to those in the more common spinful TLL realizations. Most demonstrations focus on tunnelling spectroscopy by either electron transport or local probe spectroscopy. The smoking gun signature for a TLL is thereby the presence of a pseudogap in the measured tunneling density of states at the Fermi level that, according to Eq. (17), scales as a power-law, universally, in bias voltage and temperature. Curiously, quantum Hall liquids at bulk filling factor $\nu>1$ have been reported consistent with Fermi liquid behavior[72]. Moreover, for $\nu<1$, the universal tunneling exponent was not related to the bulk filling factor[31]. While several experimental reports on 2D TIs have shown the presence of a pseudogap features within the DOS of the topological edge states at the Fermi level [149, 45, 130, 36, 146, 84, 128], consistent with a Luttinger liquid, only few examples exist [107, 146, 84] in which the universal scaling dependence has been systematically investigated. The first report of universal scaling for tunneling into the helical edge states of a QSH insulator was demonstrated for inverted InAs/GaSb heterostructures, in transport spectroscopy [107]. Fig. 4 shows more recent demonstrations for the atomically-thin QSH insulators [108] bismuthene [146] and WTe2 [84], both probed in scanning tunneling microscopy (STM). Both these studies have allowed for precise extractions of the Luttinger parameters $K=0.42$ [146] and $K=0.2\sim 0.4$ [84, 128], respectively, the latter tunable via both dielectric screening [84] and field-effect doping [128].
@@ V.3 Inter-edge Tunneling Device
Inter-edge tunneling in Luttinger liquids has gathered significant attention for its role in edge state dynamics in quantum Hall systems. Inter-edge tunneling refers to the process where electrons tunnel between different edge states, which can profoundly affect the transport properties of the system. This phenomenon is particularly interesting in fractional quantum Hall states, where the tunneling can reveal information about the fractional charge and anyonic statistics of the quasiparticles involved. Theoretical studies, such as those by Wen[168] and Kane and Fisher[87], have shown that the tunneling current in these systems follows a power-law behavior as a function of temperature and bias voltage, with exponents that depend on the Luttinger liquid parameters. These insights have been pivotal in understanding the non-trivial correlation effects and the emergent exotic physics in low-dimensional quantum systems .
More recent experimental approaches have considered devices enabling scattering and inter-edge tunneling [77] of edge modes in narrow channels [126] or quantum point contact (QPC) constrictions [163, 145, 35]. Other proposals to measure the interaction strength in helical liquids include microwave impedance measurements [61]. Examples of devices enabling inter-edge tunneling are provided in Figure5.
Figure 5: Realizations of inter-edge tunneling devices to probe TLLs in quantum Hall different 2D TIs. a-b, Tunneling spectroscopy of a graphene point contact connecting integer ($\nu=1$) and fractional ($\nu=1/3$) quantum Hall states. Panel b shows the finite bias spectroscopy of the QPC conductance and its characteristic zero-boas suppression. c-d, Point contact constriction, lithographically defined in the quantum spin Hall regime of an inverted HgTe/CdTe heterostructure. Panels a-b were taken from Ref. [35]. Panels c-d were taken from Ref. [145].
@@ VI.1 Pinning transition in 1D Bose gases
The Tomonaga-Luttinger Liquid framework describes not only 1D fermionic gases but it also apply to the low energy description of 1D bosonic gases. One-dimensional Bose gases are obtained in cold atoms setups, using either a 2-dimensional deep optical lattice that confines the atoms in each tube of the lattice or a micro-magnetic trap to freeze out the transverse degree of freedom. In many experiments, the gas is free along the longitudinal direction up to a slowly varying longitudinal potential, whose effect can be captured within a local density approximation. Thus, the system is described locally by a Galilean invariant system, in which case the Luttinger liquid parameter is linked to the speed of sound and density according to $K=\pi\rho_{0}/v$. Moreover, in most experiments the interactions between Bosons is well modeled by a contact repulsive two-body interaction of strength $g$, whose relative strength is quantified by the dimensionless parameter $\gamma=mg/(\hbar^{2}\rho_{0})$. The system then realizes the famous Lieb-Liniger model and $K$ can be computed exactly as a function of $\gamma$: it goes to 1 as $\gamma$ goes to infinity, which corresponds to the hard-core Boson limit (Tonks gas), and it takes the asymptotic value $K\simeq\pi/\sqrt{\gamma}$ for $\gamma\rightarrow 0$. The realization of a Tonks-Girardeau gas, as demonstrated in the seminal works by Weiss et al.[96] and Bloch et al.[122], has provided a profound experimental platform for exploring one-dimensional quantum systems, while Schmiedmayer et al.[74] further advanced the field by observing finite-temperature correlation functions of a Luttinger liquid through interference experiments, offering critical insights even in regimes with very high Luttinger parameters, i.e. $\gamma\rightarrow 0$. Long-wavelength physics, obtained ignoring the oscillating terms in Eq.2, reduces to the physics of phonons, or sound waves, and many experimental results obtained in 1D Bose gases can be understood within this framework, both at equilibrium and out-of-equilibrium.
@@
On the other hand, oscillating terms in Eq. 2 play an important role in the presence of an external potential whose wavelength is close to the inter-particle distance $1/\rho_{0}$. The Hamiltonian then take the form of a Sine-Gordon Hamiltonian. This model presents a quantum phase transition from a delocalized state characterized by algebraically decaying correlation functions to a localized state with a first order equal-time correlation function decaying exponentially. In the case of a periodic potential whose wavelength is exactly $1/\rho_{0}$, this transition occurs when the periodic potential amplitude exceeds a critical value $V_{c}$, which depends on $K$. Remarkably, $V_{c}$ vanishes when approaches $K_{c}$: for $K<K_{c}$ a periodic potential of vanishing amplitude is sufficient to localize the system. This transition is of Kosterlitz-Thouless type, the 2D physics arising here thanks to the equivalence between a quantum field 1D problem and a 2D thermal classical field, and the Luttinger parameter at the transition is $K_{c}=2$. Those predictions have been verified in cold atoms experiments [69, 18].

### v2

# Platforms for the realization and characterization of Tomonaga Luttinger liquids

@@ II.1 Hamiltonian and observables of the Tomonaga-Luttinger liquid
where $\phi$ and $\theta$ are bosonic fields representing collective excitations of the system — specifically, long-wavelength density and phase fluctuations. They obey the commutation relation $[\phi(x),\theta(x^{\prime})]=i\pi\Theta_{H}(x^{\prime}-x)$, $\Theta_{H}$ being the Heaviside step function, while $u$ is the velocity of low-energy excitations and $K$ is the dimensionless Tomonaga-Luttinger liquid parameter. The fermion density is expressed in terms of the operator $\phi(x)$ as
@@
for the boson annihilation operator, while (2) gives the boson density. More generally, bosons with repulsive interactions are in a Tomonaga-Luttinger liquid phase[71]. Independently of particle statistics, the charge stiffness[109] is $\mathcal{D}=uK$ while the compressibility is $\chi=K/(\pi\rho_{0}^{2}u)$, and at low temperature, the specific heat is $C_{v}=\frac{\pi k_{B}^{2}T}{3u}$. These relations allow a non-perturbative definition of $u$ and $K$. Moreover, since hard core bosons can be mapped to spin-1/2 operators[122], with $S_{j}^{+}\sim(-)^{j}\psi_{B}^{\dagger}(ja)$ and $S_{j}^{z}=\rho(x)-1/(2a)$, XXZ spin-1/2 chains also host a Tomonaga-Luttinger liquid phase.[117, 70] A fully polarized spin state maps to empty or full Jordan-Wigner fermion band, and a state with partial magnetization corresponds to the partial filling of the fermion band and is a TLL. The role of the compressibility is played by the magnetic susceptibility, $\chi=(g\mu_{B})^{2}K/(\pi u)$ where $\mu_{B}$ is the Bohr magneton and $g$ the $g$-factor, while charge stiffness is replaced by spin stiffness[152]. The magnetic heat capacity $C_{m}=\pi k_{B}^{2}T/(3u)$ has a linear temperature dependence, and the Wilson ratio, $R_{W}=\frac{4}{3}\left(\frac{\pi k_{B}}{g\mu_{B}}\right)^{2}\frac{\chi T}{C_{m}}=4K$ measures the Tomonaga-Luttinger exponent[126]. In the case of spin-1/2 fermions, the two-component Tomonaga-Luttinger liquid Hamiltonian reads
@@
with $\psi_{\sigma}(x)=\sum_{r}e^{ir\pi\rho_{0}x/2}\psi_{r,\sigma}(x)$. At half filling, $\rho_{0}=1/a$ replacing the operator $\phi_{c}$ with 00 in the spin density [] Eq.(8)] gives the bosonized expression of the spin operators in the spin-1/2 Heisenberg chain. In the helical liquid[184, 187], spin and Fermi momentum are locked together, making the Tomonaga-Luttinger liquid single component [Eq. (1)] as in the case of spinless fermions. The fermion annihilation operators reduce to
@@ II.2 Ground state correlation functions
where $\alpha$ is a short distance cutoff. The power-law singularities in Eq. (14) are determined by $n,m$ and the Tomonaga-Luttinger liquid exponent $K$. From the correlation functions above, one finds the generalized susceptibilities $\chi_{n,m}(q,\omega)$ at frequency $\omega$ and wavevector $q$ using Fourier transformation. They present power law singularities[61]
$\chi_{n,m}(q,\omega)$ $\sim$ $\left[\frac{(\omega-uq)\alpha}{u}\right]^{\frac{1}{4}\left(\frac{n}{\sqrt{K}}+m\sqrt{K}\right)^{2}-1}$ (15)
$\times$ $\left[\frac{(\omega+uq)\alpha}{u}\right]^{\frac{1}{4}\left(\frac{n}{\sqrt{K}}-m\sqrt{K}\right)^{2}-1},$
@@ III Itinerant condensed matter TLLs
Itinerant condensed matter systems, such as Tomonaga-Luttinger liquids (TLLs), represent a fascinating class of one-dimensional quantum systems characterized by their distinct properties from traditional Fermi liquids. Unlike conventional three-dimensional metals, where electron interactions can often be treated as perturbative, TLLs exhibit strong correlations that fundamentally alter their behavior. The theoretical framework of TLL provides a robust description of these phenomena. It applies not only to quantum wires[161, 5, 91, 154, 41] and carbon nanotubes[18, 195] but also to edge states[182] in the fractional quantum Hall effect and certain quasi-one-dimensional[179, 88] conductors, highlighting the universal nature of these one-dimensional systems in condensed matter physics. One specific example of itinerant condensed matter TLL is the 1D array of Josephson junctions.
@@ III.1 Josephson junction chains
In the absence of quantum fluctuations, the system would be analogous to the classical pinning of an elastic chain.[58] Such a chain is characterized by a localization length known as the Larkin length, $N_{\mathrm{L}}$. Classically, one would need to apply a force, here an external voltage, to initiate a flow of current. For chains larger than $N_{\mathrm{L}}$, the pinning force is the critical voltage divided by the number of junctions in the chain, and has dependence, $eV_{\mathrm{c}}/N\propto N_{\mathrm{L}}^{-2}\propto E_{\mathrm{0}}^{-1/3}W^{4/3}$, where $E_{\mathrm{0}}$ is the elastic energy, and W is the width of the disorder, equal to the Bloch bandwidth of the junction energy band for maximal disorder. We note $W\sim\hbar\omega_{p}$, a characteristic energy of the junctions. In order to experimentally compare chain families of widely varying $\hbar\omega_{p}$, and chain length $N$, we express the critical voltage and Bloch bandwidth as dimensionless variables, $v\equiv eV_{\mathrm{c}}/N\hbar\omega_{p}$, and $w\equiv W/\hbar\omega_{p}$. In the classical limit, $v=aw^{4/3}$, with prefactor $a=b\left(K/\Lambda\right)^{1/3}$, $b$ a constant $\mathcal{O}\left(1\right)$, and $\Lambda$ the electrostatic screening length.
@@ V Edge states and hinge states of topological matter
A natural realization of 1D electronic structure arises within the boundary modes of 2D topological insulators (TIs) (Fig. 3). Different from conventional electrical insulators (or semiconductor) in which conduction and valence bands are strictly separated by the electronic band gap, topological insulators possess a band-gap only in the bulk of the crystal while metallic boundary states persist on their surfaces and edges.
Regardless of dimensionality, what makes topological boundary modes fundamentally different from their trivial counterparts (such as e.g. Shockley [153] or Tamm[159] states) is that (i) their existence is ensured by topological invariants of the bulk band structure – a principle called bulk-boundary correspondence [75] – and that (ii) these systems often carry (at least approximately) linearly dispersing 1D Dirac fermions at their boundaries that are either chiral or helical, spin-degenerate or spinless, depending on which class of topological insulator is being considered (Fig. 3d).
@@
The first-known 2D topological phase of matter is the quantum Hall effect (QHE) [107], later put into the context of band topology by Thouless [166]. In the QHE, a strong magnetic field $H$ externally applied perpendicular to the plane of a 2D electron gas forces electrons into closed orbits leading to the opening of a gap in the 2D bulk density of states (Landau quantization). Gapless chiral edge states[181] persist at the boundary, spin-degenerate at low magnetic field (Fig. 3b).
@@ V.2 Experimental demonstrations of chiral and helical TLLs
A significant amount of experimental work has focused on investigating Luttinger liquids within the chiral edges of quantum Hall states (see Refs.[33, 57] for comprehensive reviews). Indeed, quantum Hall states offer a considerable degree of tunability [57], depending on the electronic filling factor $\nu$ determining the number of chiral edge states. Possible scenarios include co-propagating pairs of edge modes with opposite spin ($\nu=2$), spin-polarized single modes ($\nu=1$), fractional modes, and even counter-propagating (spin-degenerate or non-degenerate) modes of adjacent quantum Hall domains separated by a thin gate-electrode [94, 136, 57]. TLLs have just recently been detected in the helical edge states of QSH insulators [115, 157, 89, 156], while no Luttinger liquid claims have been made, to our knowledge, for the chiral edges of QAH (or Chern) insulators [190, 35].
@@
Similar to early investigations in 1D wires[41] and quantum Hall edge states [34, 33], experimental probes of chiral and helical Luttinger liquids in 2D TIs have largely focused on tunnelling spectroscopy by either electron transport or local probe spectroscopy. The smoking gun signature for a TLL is thereby the presence of a pseudogap in the measured tunneling density of states at the Fermi level that, according to Eq. (17), scales as a power-law, universally, in bias voltage and temperature. Curiously, quantum Hall liquids at bulk filling factor $\nu>1$ have been reported consistent with Fermi liquid behavior[77]. Moreover, for $\nu<1$, the universal tunneling exponent was not related to the bulk filling factor[33]. While several experimental reports on 2D TIs have shown the presence of a pseudogap features within the DOS of the topological edge states at the Fermi level [160, 49, 140, 39, 157, 89, 138], consistent with a Luttinger liquid, only few examples exist [115, 157, 89] in which the universal scaling dependence has been systematically investigated. The first report of universal scaling for tunneling into the helical edge states of a QSH insulator was demonstrated for inverted InAs/GaSb heterostructures, in transport spectroscopy [115]. Fig. 4 shows more recent demonstrations for the atomically-thin QSH insulators [116] bismuthene [157] and WTe2 [89], both probed in scanning tunneling microscopy (STM). Both these studies have allowed for precise extractions of the Luttinger parameters $K=0.42$ [157] and $K=0.2\sim 0.4$ [89, 138], respectively, the latter tunable via both dielectric screening [89] and field-effect doping [138].
@@ V.3 Inter-edge Tunneling Device
Inter-edge tunneling in Luttinger liquids has gathered significant attention for its role in edge state dynamics in quantum Hall systems [57]. Inter-edge tunneling refers to the process where electrons tunnel between different edge states, which can profoundly affect the transport properties of the system. This phenomenon is particularly interesting in fractional quantum Hall states, where the tunneling can reveal information about the fractional charge and anyonic statistics of the quasiparticles involved. Theoretical studies, such as those by Wen[183] and Kane and Fisher[95], have shown that the tunneling current in these systems follows a power-law behavior as a function of temperature and bias voltage, with exponents that depend on the Luttinger liquid parameters. These insights have been pivotal in understanding the non-trivial correlation effects and the emergent exotic physics in low-dimensional quantum systems. More recent experimental approaches have considered devices enabling scattering and inter-edge tunneling [82] of edge modes in narrow channels [136] or quantum point contact (QPC) constrictions [178, 38].
While the bulk of experimental demonstrations to date have focused on inter-edge tunneling devices in the (integer and fractional) quantum Hall regimes [57, 178, 38], recent work also includes point contact spectroscopy measurements of helical QSH edges [156]. Other proposals to measure the interaction strength in helical liquids include microwave impedance measurements [66]. Examples of devices enabling inter-edge tunneling are provided in Figure5.
Figure 5: Realizations of inter-edge tunneling devices to probe TLLs at the edges of quantum Hall and quantum spin Hall insulators. a-b, Tunneling spectroscopy of a graphene point contact connecting integer ($\nu=1$) and fractional ($\nu=1/3$) quantum Hall states. Panel b shows the finite bias spectroscopy of the QPC conductance and its characteristic zero-boas suppression. c-d, Point contact constriction, lithographically defined in the quantum spin Hall regime of an inverted HgTe/CdTe heterostructure. Panels a-b were taken from Ref. [38]. Panels c-d were taken from Ref. [156].
@@ VI.1 Pinning transition in 1D Bose gases
The Tomonaga-Luttinger Liquid framework describes not only 1D fermionic gases but it also applies to the low energy description of 1D bosonic gases. One-dimensional Bose gases are obtained in cold atoms setups, using either a 2-dimensional deep optical lattice that confines the atoms in each tube of the lattice or a micro-magnetic trap to freeze out the transverse degree of freedom. In many experiments, the gas is free along the longitudinal direction up to a slowly varying longitudinal potential, whose effect can be captured within a local density approximation. Thus, the system is described locally by a Galilean invariant system, in which case the Luttinger liquid parameter is linked to the speed of sound and density according to $K=\pi\rho_{0}/v$. Moreover, in most experiments the interactions between Bosons is well modeled by a contact repulsive two-body interaction of strength $g$, whose relative strength is quantified by the dimensionless parameter $\gamma=mg/(\hbar^{2}\rho_{0})$. The system then realizes the famous Lieb-Liniger model and $K$ can be computed exactly as a function of $\gamma$: it goes to 1 as $\gamma$ goes to infinity, which corresponds to the hard-core Boson limit (Tonks gas), and it takes the asymptotic value $K\simeq\pi/\sqrt{\gamma}$ for $\gamma\rightarrow 0$. The realization of a Tonks-Girardeau gas, as demonstrated in the seminal works by Weiss et al.[104] and Bloch et al.[132], has provided a profound experimental platform for exploring one-dimensional quantum systems, while Schmiedmayer et al.[79] further advanced the field by observing finite-temperature correlation functions of a Luttinger liquid through interference experiments, offering critical insights even in regimes with very high Luttinger parameters, i.e. $\gamma\rightarrow 0$. Long-wavelength physics, obtained ignoring the oscillating terms in Eq.2, reduces to the physics of phonons, or sound waves, and many experimental results obtained in 1D Bose gases can be understood within this framework, both at equilibrium and out-of-equilibrium.
@@
On the other hand, oscillating terms in Eq. 2 play an important role in the presence of an external potential whose wavelength is close to the inter-particle distance $1/\rho_{0}$. The Hamiltonian then takes the form of a Sine-Gordon Hamiltonian. This model presents a quantum phase transition from a delocalized state characterized by algebraically decaying correlation functions to a localized state with a first order equal-time correlation function decaying exponentially. In the case of a periodic potential whose wavelength is exactly $1/\rho_{0}$, this transition occurs when the periodic potential amplitude exceeds a critical value $V_{c}$, which depends on $K$. Remarkably, $V_{c}$ vanishes when approaches $K_{c}$: for $K<K_{c}$ a periodic potential of vanishing amplitude is sufficient to localize the system. This transition is of Kosterlitz-Thouless type, the 2D physics arising here thanks to the equivalence between a quantum field 1D problem and a 2D thermal classical field, and the Luttinger parameter at the transition is $K_{c}=2$. Those predictions have been verified in cold atoms experiments [74, 20].